# GazeGaussian: High-Fidelity Gaze Redirection with 3D Gaussian Splatting

### v1

@@ Introduction
Finally, we apply rigid rotations and translations to transform Gaussians in the canonical space to the world space. Then, these Gaussians are rasterized into the feature maps. The above face-only branch can be formulated as:
$\begin{split}\mathcal{M}_{f}&=\mathcal{R}(\{X^{f},C^{f},Q^{f},S^{f},A^{f}\})\\ &=\mathcal{R}(\boldsymbol{\Phi}^{f}(\boldsymbol{X}_{0}^{f},\boldsymbol{F}_{0}^{f},\boldsymbol{Q}_{0}^{f},\boldsymbol{S}_{0}^{f},\boldsymbol{A}_{0}^{f};\theta,\beta)),\end{split}$ (5)
@@
For the eye branch, we also construct canonical neutral eye Gaussians with attributes $\{\boldsymbol{X}_{0}^{e},\boldsymbol{F}_{0}^{e},\boldsymbol{Q}_{0}^{e},\boldsymbol{S}_{0}^{e},\boldsymbol{A}_{0}^{e}\}$. These attributes share the same dimensionality as those in the face-only branch, except that $\boldsymbol{S}_{0}^{e}\in\mathbb{R}^{N\times 1}$ is constrained to be spherical, aligning with the rotational properties of the eyeball. Next, we describe the process of applying offsets to each Gaussian attribute.
@@
Positions $X^{e}$ of the Gaussians. Directly applying the same deformation strategy as for the face branch would fail to fully leverage the unique characteristics of eyeball rotational motion, resulting in insufficient gaze redirection accuracy. Therefore, we first rotate the eye Gaussians in the canonical space and then incorporate the eye geometry information from the latent codes of different subjects to generate biases. Since the gaze labels may contain noise, directly using the normalized gaze direction $\varphi$ to rotate the Gaussians would lead to numerical optimization errors. To address this, we optimize two separate MLPs: ${}_{rot}^{gaze}\mathcal{F}^{e}\in\boldsymbol{\Phi}^{e}$ and ${}_{def}^{exp}\mathcal{F}^{e}\in\boldsymbol{\Phi}^{e}$ to predict the biases for Gaussian rotation and displacement.
$\begin{split}X^{e}=_{def}^{exp}\mathcal{F}^{f}(\boldsymbol{X}_{0}^{e},\theta)+_{rot}^{gaze}\mathcal{F}^{e}(\boldsymbol{X}_{0}^{e},\varphi)\boldsymbol{X}_{0}^{e},\end{split}$ (6)
Since eyes are relatively small and mainly influenced by the gaze direction, $\lambda$ used in the face is omitted here.
Color $C^{e}$ of the Gaussians. The color of the eye region is influenced by the gaze direction and latent codes. We use two MLPs: ${}_{col}^{exp}\mathcal{F}^{e}\in\boldsymbol{\Phi}^{e}$ and ${}_{col}^{gaze}\mathcal{F}^{e}\in\boldsymbol{\Phi}^{e}$ to predict it:
$C^{e}=_{att}^{exp}\mathcal{F}^{e}(\boldsymbol{F}_{0}^{e},\theta)\\ +_{col}^{gaze}\mathcal{F}^{e}(\boldsymbol{X}_{0}^{e},\varphi),$ (7)
Rotation, Scale and Opacity $\{Q^{e},S^{e},A^{e}\}$ of the Gaussians. We just use another two attribute MLPs ${}_{att}^{exp}\mathcal{F}^{e}\in\boldsymbol{\Phi}^{e}$ and ${}_{att}^{gaze}\mathcal{F}^{e}\in\boldsymbol{\Phi}^{e}$ to predict their shift.
$\begin{split}\{Q^{e},S^{e},A^{e}\}=\{\mathbf{Q}_{0}^{e},\mathbf{S}_{0}^{e},\mathbf{A}_{0}^{e}\}&+_{att}^{exp}\mathcal{F}^{e}(\boldsymbol{F}_{0}^{e},\theta)\\ &+_{att}^{gaze}\mathcal{F}^{e}(\boldsymbol{F}_{0}^{e},\varphi),\end{split}$ (8)
Finally, we transform Gaussians in the canonical space to the world space. Then these eye Gaussians are rasterized into the feature maps. The eye branch is formulated as:
$\begin{split}\mathcal{M}_{e}&=\mathcal{R}(\{X^{e},C^{e},Q^{e},S^{e},A^{e}\})\\ &=\mathcal{R}(\boldsymbol{\Phi}^{e}(\boldsymbol{X}_{0}^{e},\boldsymbol{F}_{0}^{e},\boldsymbol{Q}_{0}^{e},\boldsymbol{S}_{0}^{e},\boldsymbol{A}_{0}^{e};\theta,\varphi)),\end{split}$ (9)
@@
To obtain the full head rendering, we simply concat the two-stream Gaussians and rasterized them into feature maps:
$\mathcal{M}_{h}$ =($=\mathcal{R}(\{X^{f},C^{f},Q^{f},S^{f},A^{f}\}$
$\{X^{e},C^{e},Q^{e},S^{e},A^{e}\}),$ (10)
Expression-Guided Neural Renderer
@@
After obtaining the rasterized feature maps from Gaussians, a UNet-like neural renderer $\mathbf{R}$ opts to synthesize the final face-only, eyes, and head images $\{\mathcal{I}_{f},\mathcal{I}_{e},\mathcal{I}_{h}\}$:
$\{\mathcal{I}_{f},\mathcal{I}_{e},\mathcal{I}_{h}\}=\mathbf{R}(\{\mathcal{M}_{f},\mathcal{M}_{e},\mathcal{M}_{h}\}),$ (11)
@@
To enhance the generalization ability across different subjects, we inject the latent codes $\theta$ into the neural renderer through a slice cross-attention module. Let $\boldsymbol{F}_{b}$ represent the bottleneck feature obtained from the encoder of $\mathbf{R}$. We utilize the latent codes to query this bottleneck feature, using it as a conditional signal to guide the renderer’s synthesis process. The guiding process can be formulated as:
$\boldsymbol{F}^{\prime}_{b}=\boldsymbol{F}_{b}+\boldsymbol{F}_{b}\cdot\text{Attn}(q=\theta,k=\boldsymbol{F}_{b},v=\boldsymbol{F}_{b})),$ (12)
@@
where $\text{Attn}(\cdot)$ denotes the cross-attention operation that fuses the latent codes with the bottleneck feature. Then the refined feature $\boldsymbol{F}^{\prime}_{b}$ is decoded as final images.
Training
@@
GazeGaussian Initialization. Initialization for the 3D Gaussians (3DGS) is crucial for stable optimization. Following Gaussian Head Avatar, we initialize the two-stream Gaussians using the neutral mesh extracted from an SDF field. This neutral mesh provides a coarse geometry and texture, which are used to initialize the positions and features of the Gaussians. To decouple the face-only and eye regions, we compute the 3D neutral landmarks and use learnable parameters to define the vertices near the eyes as the initial Gaussians for the eye region, while the rest of the head is used to initialize the face-only Gaussians. Additionally, we transfer the parameters of all deformation and color MLPs while the MLPs for attribute prediction and the expression-guided neural renderer are randomly initialized.
@@
Image Synthesis Loss. The masked ground truth image $\mathcal{I}_{gt}$ is used to supervise the rendered images ${\mathcal{I}_{f},\mathcal{I}_{e},\mathcal{I}_{h}}$, corresponding to the face-only, eyes, and head regions, respectively. Additionally, we enforce the first three channels of the feature maps ${\mathcal{M}_{f},\mathcal{M}_{e},\mathcal{M}_{h}}$ to learn the RGB colors. For each rendered image and its corresponding feature map, we apply the same loss functions. Taking the rendered eye image as an example, we mask the ground truth image using an eye mask and then apply L1 loss, SSIM loss, and LPIPS loss on the masked image:
@@
where $\lambda_{SSIM}=\lambda_{VGG}=0.1$ is the weight of loss. The image synthesis loss is the sum of the three renderer images and three feature maps:
@@
Gaze Redirection Loss. To improve task-specific performance and eliminate task-relevant inconsistencies between the target image $\mathcal{I}_{gt}$ and the reconstructed head image $\mathcal{I}_{h}$, we adopt the functional loss used in STED (Zheng et al. 2020) and GazeNeRF (Ruzzi et al. 2023). The gaze redirection loss can be formulated as:
$\begin{split}\mathcal{L}_{\mathcal{G}}(\mathcal{I}_{h},\mathcal{I}_{gt})&=\mathcal{E}_{\text{ang}}(\psi^{g}(\mathcal{I}_{wf}),\psi^{g}(\mathcal{I}_{gt}))\\ \mathcal{E}_{\text{ang}}(\mathbf{v},\ \hat{\mathbf{v}})&=\arccos\frac{\mathbf{v}\cdot{\hat{\mathbf{v}}}}{\|\mathbf{v}\|\,\|{\hat{\mathbf{v}}}\|}\,,\end{split}$ (15)
Experiments
To demonstrate the effectiveness of GazeGaussian, we first conduct a within-dataset comparison on the ETH-XGaze dataset (Zhang et al. 2020), testing GazeGaussian alongside state-of-the-art gaze redirection and head generation methods. Next, we perform a cross-dataset comparison on ColumbiaGaze (Smith et al. 2013), MPIIFaceGaze (Zhang et al. 2017, 2015), and GazeCapture (Krafka et al. 2016) to assess generalization. We also conduct an ablation study to analyze the contributions of each component in GazeGaussian. Additionally, we validate the impact of synthesized data on gaze estimator performance in the supplementary materials. Due to space limitations, please refer to the supplementary for more details on the experiment and visualization results.
@@
Figure 3: Within-dataset comparison: Visualization of generated images from the ETH-XGaze test set using our GazeGaussian, GazeNeRF, and Gaussian Head Avatar. All faces are masked to remove the background. GazeGaussian generates photo-realistic images with the target gaze direction, preserving identity and facial details. In contrast, GazeNeRF loses identity information and facial details, while Gaussian Head Avatar fails to manipulate the gaze direction effectively.
Experimental Settings
@@
Dataset Pre-processing. Following GazeNeRF’s preprocessing, we normalize raw images (Zhang, Sugano, and Bulling 2018; Sugano, Matsushita, and Sato 2014) and resize them into a resolution 512$\times$512. To enable separate rendering of the face and eyes regions, we generate masks using face parsing models (zllrunning 2019). We also use the 3D face tracking method from (Xu et al. 2024) to produce identity and expression codes and camera poses for the input of our method. For consistency, gaze labels are converted to pitch-yaw angles in the head coordinate system across all datasets. Details are provided in the supplementary materials.
Baselines.
@@
We compare our method with the self-supervised gaze redirection approach STED (Zheng et al. 2020), along with NeRF-based models such as HeadNeRF (Hong et al. 2022b) and the state-of-the-art method GazeNeRF (Ruzzi et al. 2023), as well as the latest 3DGS-based head synthesis method, Gaussian Head Avatar (Xu et al. 2024). As the NeRF-based methods, NeRF-Gaze (Yin et al. 2024) and Wang et al. (Wang et al. 2023) are not yet open-sourced, they are not available for inclusion in our comparisons.
Metrics.
@@
We evaluate all models using four categories: redirection accuracy, image quality, identity preservation, and rendering speed. Redirection accuracy is measured by gaze and head poses angular errors, using a ResNet50 (He et al. 2016)-based estimator, as in GazeNeRF (Ruzzi et al. 2023). Image quality is assessed with SSIM, PSNR, LPIPS, and FID. Identity preservation is evaluated with FaceX-Zoo (Wang et al. 2021a), comparing identity consistency between redirected and ground-truth images. Rendering speed is reported as average FPS.
@@
Figure 4: Cross-dataset comparison: Visualization of generated images from the MPIIFaceGaze test set using our GazeGaussian, GazeNeRF, and Gaussian Head Avatar. Please refer to the supplementary for more visualization.
Within-dataset Comparison
Following the experimental setup of GazeNeRF, we perform a within-dataset evaluation to compare the performance of GazeGaussian with other state-of-the-art methods. All models are trained using 14.4K images derived from 10 frames per subject, with 18 camera view images per frame, covering 80 subjects in the ETH-XGaze training set. The evaluation is conducted on the person-specific test set of the ETH-XGaze dataset. This test set consists of 15 subjects, each with 200 images annotated with gaze and head pose labels. We follow the pairing setting in GazeNeRF, which pairs these 200 labeled images per subject as input and target samples, and the same pairings are used across all models to ensure fairness.
Tab. 1 presents the quantitative results of GazeGaussian alongside baseline methods. It can be observed that GazeGaussian consistently outperforms prior methods across all metrics. Specifically, our approach achieves the lowest errors in both gaze and head redirection ($6.622^{\circ}$ and $2.128^{\circ}$, respectively), demonstrating its superior precision in gaze and head control. Compared to the previous SOTA method GazeNeRF, which applies rotation to feature map for gaze redirection, GazeGaussian adopts a Gaussian eye rotation to explicitly control eye movement. Such a technique not only improves redirection accuracy but also significantly boosts rendering quality. Additionally, GazeGaussian achieves a rendering speed of 74 FPS, nearly doubling the performance of GazeNeRF, underscoring its efficiency. In contrast, Gaussian Head Avatar (GHA), the latest model built on Gaussian-based representations, struggles to deliver competitive performance in gaze and head redirection tasks. The lack of dedicated mechanisms for gaze disentanglement and explicit eye region modeling in GHA leads to poor performance. By decoupling the face and eye representation with two-stream Gaussians, GazeGaussian offers both higher accuracy and better visual quality, particularly in challenging scenarios involving extreme head poses or subtle gaze variations.
We present a qualitative comparison of different methods in Fig. 3. GHA struggles to preserve personal identity in the generated face images, which is quantitatively verified as the low ‘identity similarity’ in Tab. 1. Moreover, GHA produces blurred and unrealistic eye regions, significantly degrading the visual quality of gaze redirection. GazeNeRF, which implicitly rotates the feature map, fails to effectively control eye appearance under extreme gaze directions (as shown in the last row). Furthermore, it struggles with rendering fine-grained facial details and exhibits notable artifacts in hair rendering, particularly in the last two rows. Overall, the inability to accurately handle eye details in both GHA and GazeNeRF limits their effectiveness in gaze redirection. In contrast, GazeGaussian consistently produces highly realistic results, even under challenging conditions, setting a new benchmark for gaze redirection tasks.
Table 3: Component-wise ablation study of GazeGaussian on the ETH-XGaze dataset in terms of gaze and head redirection errors in degree, redirection image quality (SSIM, PSNR, LPIPS and FID), and identity similarity.
@@
Cross-dataset Comparison
To access the generalization capability of GazeGaussian , we perform a cross-dataset evaluation on three other datasets: ColumbiaGaze, MPIIFaceGaze, and the test set of GazeCapture. The training setup remains consistent with the within-dataset evaluation, using the same model configurations and trained parameters.
The results shown in Tab. 2 and Fig. 4 demonstrate that GazeGaussian consistently outperforms all other methods across the three datasets and all evaluation metrics. By introducing a novel expression-guided neural renderer, GazeGaussian can retain facial details across various subjects. On the other hand, GHA’s performance is limited by its modeling strategy, showing poor adaptability to unseen datasets. It produces less clear eye regions and achieves significantly lower identity similarity scores compared to GazeGaussian . These results further validate the superiority of GazeGaussian , making it a more robust choice for handling diverse datasets and complex gaze redirection tasks. Please refer to supplementary material for more visualization results on the cross-dataset evaluation.

### v2

@@ 4.2.1 Face Deformation
where $\kappa^{f}=\{R^{f},S^{f},\alpha^{f}\}$ and $\kappa_{0}^{f}=\{R_{0}^{f},S_{0}^{f},\alpha_{0}^{f}\}$. Finally, we apply rigid rotations and translations to transform Gaussians in the canonical space to the world space. Then, these Gaussians are rasterized into the feature maps:
$\mathcal{M}_{f}=\mathcal{R}(\{\mu^{f},c^{f},R^{f},S^{f},\alpha^{f}\}),$ (3)
@@ 4.2.2 Eye Rotation
Different from Gaussian Head Avatar [51], we design an eye branch for accurate eye modeling and gaze redirection. For the eye branch, we construct canonical neutral eye Gaussians with attributes $\{\mu_{0}^{e},z_{0}^{e},R_{0}^{e},S_{0}^{e},\alpha_{0}^{e}\}$. These attributes share the same dimension as those in the face-only branch, except that $S_{0}^{e}\in\mathbb{R}^{N\times 1}$ is constrained to be spherical, aligning with the rotational properties of the eyeball. Since eyes are relatively small and mainly influenced by the gaze direction, $\lambda$ used in the face is omitted here.
@@
Directly applying the same face deformation strategy fails to fully leverage the unique characteristics of eyeball rotational motion, resulting in inaccurate gaze redirection. Therefore, we first rotate the eye Gaussians in the canonical space and then incorporate the expressive latent codes $\tau$ to generate deformation offsets. Since the gaze labels contain noise, directly using the normalized gaze direction $\varphi$ to rotate the Gaussians would lead to numerical optimization errors. To address this, we optimize two separate MLPs: $E_{\mu}^{e}\in E^{e}$ and $G_{\mu}^{e}\in G^{e}$ to predict the biases for Gaussian eye rotation.
$\begin{split}\mu^{e}=E_{\mu}^{e}(\mu_{0}^{e},\tau)+G_{\mu}^{e}(\mu_{0}^{e},\varphi)\mu_{0}^{e},\end{split}$ (4)
Other attributes are deformed in a similar way with $\{E_{c}^{e},E_{a}^{e}\}\in E^{e}$ and $\{G_{c}^{e},G_{a}^{e}\}\in G^{e}$:
$\begin{split}c^{e}&=E_{c}^{e}(z_{0}^{e},\tau)+G_{c}^{e}(z_{0}^{e},\varphi),\\ \kappa^{e}&=\kappa_{0}^{e}+E_{\kappa}^{e}(z_{0}^{e},\tau)+G_{\kappa}^{e}(z_{0}^{e},\varphi),\end{split}$ (5)
where $\kappa^{e}=\{R^{e},S^{e},\alpha^{e}\}$ and $\kappa_{0}^{e}=\{R_{0}^{e},S_{0}^{e},\alpha_{0}^{e}\}$. By the proposed explicit eye rotation, GazeGaussian succeeds in face and eye disentanglement, which is ignored in the Gaussian Head Avatar [51]. Finally, eye Gaussians are transformed into the world space. Then these eye Gaussians are rasterized into the feature maps:
$\mathcal{M}_{e}=\mathcal{R}(\{\mu^{e},c^{e},R^{e},S^{e},\alpha^{e}\}),$ (6)
@@
To obtain the full head rendering, we simply concat the two-stream Gaussians $\{\mu^{h},c^{h},R^{h},S^{h},\alpha^{h}\}=\text{Concat}(\{\mu^{f},c^{f},R^{f},S^{f},\alpha^{f}\},\{\mu^{e},c^{e},R^{e},S^{e},\alpha^{e}\})$ and rasterized them into head feature maps:
$\mathcal{M}_{h}=\mathcal{R}(\{\mu^{h},c^{h},R^{h},S^{h},\alpha^{h}\}),$ (7)
4.3 Expression-Guided Neural Renderer
@@
To filter noise caused by uneven lighting or camera chromatic aberration [51], a UNet-like neural renderer $\mathcal{U}$ opts to rasterize feature maps into the final face-only, eyes, and head images $\{\mathcal{I}_{f},\mathcal{I}_{e},\mathcal{I}_{h}\}$:
$\{\mathcal{I}_{f},\mathcal{I}_{e},\mathcal{I}_{h}\}=\mathcal{U}(\{\mathcal{M}_{f},\mathcal{M}_{e},\mathcal{M}_{h}\}),$ (8)
@@
However, the vanilla renderer used in Gaussian Head Avatar [51] and GazeNeRF [36] focuses solely on image post-processing, resulting in a loss of subject-specific facial details. To further improve generalization ability by providing subject-specific information, we inject the expressive latent codes $\tau$ into the neural renderer through a slice cross-attention module. Let $z_{b}$ represent the bottleneck feature obtained from the encoder of $\mathcal{U}$. We utilize the latent codes to query this bottleneck feature, using it as a conditional signal to guide the renderer’s synthesis process. The guiding process can be formulated as:
$z^{\prime}_{b}=z_{b}+z_{b}\cdot\text{Attn}(q=\tau,k=z_{b},v=z_{b})),$ (9)
@@
where $\text{Attn}(\cdot)$ denotes the cross-attention operation that fuses the latent codes with the bottleneck feature. Then the refined feature $z^{\prime}_{b}$ is decoded as final images.
4.4 Training
@@
GazeGaussian Initialization. Initialization for the 3D Gaussians (3DGS) is crucial for stable optimization. Following Gaussian Head Avatar [51], we initialize the two-stream Gaussians using the neutral mesh extracted from an SDF field. This neutral mesh provides a coarse geometry and texture, which are used to initialize the positions and features of the Gaussians. To decouple the face-only and eye regions, we compute the 3D neutral landmarks and use learnable parameters to define the vertices near the eyes as the initial Gaussians for the eye region, while the rest of the head is used to initialize the face-only Gaussians. Additionally, we transfer the parameters of all deformation and color MLPs while the MLPs for attribute prediction and the expression-guided neural renderer are randomly initialized.
@@
Image Synthesis Loss. The masked ground truth image $\mathcal{I}_{gt}$ is used to supervise the rendered images ${\mathcal{I}_{f},\mathcal{I}_{e},\mathcal{I}_{h}}$, corresponding to the face-only, eyes, and head regions, respectively. Additionally, we enforce the first three channels of the feature maps ${\mathcal{M}_{f},\mathcal{M}_{e},\mathcal{M}_{h}}$ to learn RGB colors. For rendered images and the first three channels of feature maps, we apply the same loss functions. For brevity, we take the loss for the rendered eye image $\mathcal{I}_{e}$ as an example. We mask the ground truth image using an eye mask and then apply L1 loss, SSIM loss, and LPIPS loss on the masked image:
@@
where $\lambda_{SSIM}=\lambda_{VGG}=0.1$ is the weight of loss. The image synthesis loss is the sum of the three rendered images and three feature maps:
@@
Gaze Redirection Loss. To improve task-specific performance and eliminate task-relevant inconsistencies between the target image $\mathcal{I}_{gt}$ and the reconstructed head image $\mathcal{I}_{h}$, we adopt the functional loss used in STED [59] and GazeNeRF [36]. The gaze redirection loss can be formulated as:
$\begin{split}\mathcal{L}_{\mathcal{G}}(\mathcal{I}_{h},\mathcal{I}_{gt})&=\mathcal{E}_{\text{ang}}(\psi^{g}(\mathcal{I}_{h}),\psi^{g}(\mathcal{I}_{gt})),\\ \mathcal{E}_{\text{ang}}(\varphi,\ \hat{\varphi})&=\arccos\frac{\varphi\cdot{\hat{\varphi}}}{\|\varphi\|\,\|{\hat{\varphi}}\|}\,,\end{split}$ (12)
@@
Figure 3: Within-dataset visualization: head images are generated from the ETH-XGaze test set using our GazeGaussian, GazeNeRF, and Gaussian Head Avatar. GazeGaussian generates photo-realistic images with the target gaze direction, preserving identity and facial details. In contrast, GazeNeRF loses identity information and facial details, while Gaussian Head Avatar fails to manipulate the gaze direction.
Figure 4: Cross-dataset visualization: head images are from the MPIIFaceGaze test set using our GazeGaussian, GazeNeRF, and Gaussian Head Avatar. Please refer to Sec. 12.4 in the supplementary for more visualization.
5 Experiments
To demonstrate the effectiveness of GazeGaussian, we first conduct a within-dataset comparison on the ETH-XGaze dataset [58], testing GazeGaussian alongside state-of-the-art gaze redirection and head generation methods. Next, we perform a cross-dataset comparison on ColumbiaGaze [39], MPIIFaceGaze [56, 55], and GazeCapture [21] to assess generalization. We also conduct an ablation study to analyze the contributions of each component in GazeGaussian. All experimental settings strictly follow GazeNeRF [36], including dataset preprocessing, the ETH-XGaze split (14.4K images for efficient training), image pairs for evaluation, and the same pre-trained gaze estimator. Additionally, we validate the impact of synthesized data on gaze estimator performance in Sec. 5.5. Due to space limitations, please refer to Sec. 11 and Sec. 12 in the supplementary for more details on the experiment and visualization results.
5.1 Experimental Settings
@@
Dataset Pre-processing. Following preprocessing in GazeNeRF [36], we normalize raw images [57, 40] and resize them into a resolution 512$\times$512. To enable separate rendering of the face and eyes regions, we generate masks using face parsing models [63]. We also use the 3D face tracking method from [51] to produce identity and expression codes and camera poses for the input of our method. For normalization, gaze labels provided by datasets are converted to pitch-yaw angles in the head coordinate system across all datasets. More details are provided in the supplementary materials.
@@
Baselines. We compare our method with the self-supervised gaze redirection approach STED [59], along with NeRF-based models such as HeadNeRF [15] and the state-of-the-art method GazeNeRF [36], as well as the latest 3DGS-based head synthesis method, Gaussian Head Avatar [51]. As the NeRF-based methods, NeRF-Gaze [52] and Wang et al. [42] are not yet open-sourced, they are not available for inclusion in our comparisons.
@@
Metrics. We evaluate all models using four categories: redirection accuracy, image quality, identity preservation (ID), and rendering speed. Redirection accuracy is measured by gaze and head poses angular errors, using the same ResNet50 [13]-based estimator in GazeNeRF [36]. Image quality is assessed with SSIM, PSNR, LPIPS, and FID. Identity preservation (ID) is evaluated with FaceX-Zoo [43], comparing identity consistency between redirected and ground-truth images. Rendering speed is reported as average FPS.
5.2 Within-dataset Comparison
Following the experimental setup of GazeNeRF, we perform a within-dataset evaluation to compare the performance of GazeGaussian with other state-of-the-art methods. All models are trained using 14.4K images derived from 10 frames per subject, with 18 camera view images per frame, covering 80 subjects in the ETH-XGaze training set. For a fair comparison, the training set is strictly the same as GazeNeRF. Given that Gaussian Head Avatar is not specifically designed for gaze redirection and does not support gaze as a condition, only expressive latent codes are provided for its training and evaluation. The evaluation is conducted on the person-specific test set of the ETH-XGaze dataset. This test set consists of 15 subjects, each with 200 images annotated with gaze and head pose labels. We follow the pairing setting in GazeNeRF, which pairs these 200 labeled images per subject as input and target samples, and the same pairings are used across all models to ensure fairness.
Tab. 1 presents the quantitative results of GazeGaussian alongside baseline methods. It can be observed that GazeGaussian consistently outperforms prior methods across all metrics. Compared to the previous SOTA method GazeNeRF, which only applies rotation to feature maps for implicit gaze redirection, GazeGaussian adopts a Gaussian eye rotation to explicitly control eye movement. Such a technique improves redirection precision, resulting in high-quality head image synthesis. Additionally, GazeGaussian achieves a rendering speed of 74 FPS, nearly doubling the performance of GazeNeRF, underscoring its efficiency. In contrast, Gaussian Head Avatar (GHA), the latest 3DGS-based model, fails to achieve high performance in gaze redirection. The lack of dedicated mechanisms for gaze disentanglement and explicit eye region modeling in GHA leads to poor gaze redirection performance. By decoupling the face and eye representation with two-stream Gaussians, GazeGaussian offers both higher accuracy and better visual quality, particularly in challenging scenarios involving extreme head poses or subtle gaze variations. Furthermore, GHA is limited to animating a single avatar, whereas GazeGaussian with the proposed neural-guided renderer achieves better generalization ability and identity preservation across different subjects.
We present a visualization comparison of different methods in Fig. 3. GHA fails to generalize to different subjects and struggles to preserve personal identity in the rendered images, which is quantitatively verified as the low ‘identity similarity’ in Tab. 1. Moreover, GHA produces blurred and unrealistic eye regions, significantly degrading the visual quality of gaze redirection. GazeNeRF, which implicitly rotates the feature map, fails to effectively control eye appearance under extreme gaze directions (as shown in the last row). Furthermore, GazeNeRF struggles with rendering fine-grained facial details and exhibits notable artifacts in hair rendering. The failure to precisely render eye details in GHA and GazeNeRF restricts their capability in gaze redirection. In contrast, GazeGaussian consistently produces highly realistic results, even under challenging conditions, setting a new benchmark for gaze redirection tasks.
@@ 5.2 Within-dataset Comparison
Table 3: Component-wise ablation study of GazeGaussian on the ETH-XGaze dataset in terms of gaze and head redirection errors in degree, redirection image quality (SSIM, PSNR, LPIPS and FID), and identity similarity (ID).
@@
5.3 Cross-dataset Comparison
To access the generalization capability of GazeGaussian, we train all methods on the ETH-Xgaze dataset and evaluate them on three other datasets: ColumbiaGaze, MPIIFaceGaze, and the test set of GazeCapture. The training setup remains consistent with the within-dataset evaluation, using the same model configurations and trained parameters.
The results shown in Tab. 2 and Fig. 4 demonstrate that GazeGaussian consistently outperforms all other methods across the three datasets and all evaluation metrics. By introducing a novel expression-guided neural renderer, GazeGaussian gains better generalization ability by integrating person-specific features, resulting in optimal identity preservation. On the other hand, GHA’s performance is limited by its modeling strategy, showing poor adaptability to unseen datasets. It produces less clear eye regions and achieves significantly lower identity similarity scores compared to GazeGaussian. These results further validate the superiority of GazeGaussian, making it a more robust choice for handling diverse datasets and complex gaze redirection tasks.